# Uncertainty-Aware Explanations Through Probabilistic Self-Explainable Neural Networks

### v1

@@ 3.3 Capturing the explanation uncertainty
Regarding epistemic explanatory uncertainty, as explained above, a high uncertainty will be achieved for those inputs for which the distances to all the prototypes are large. On that account, the minimum sample mean: $\bar{\delta}_{x,i*}=\min_{1\leq j\leq c}\bar{\delta}_{x,j}$ represents a suitable base metric for evaluating this uncertainty. Furthermore, in order to determine the severity of the uncertainty in a normalized and scale-independent manner, a quantile-based metric will be defined, close to the approach proposed in Petrov and Kwiatkowska (2022). This metric will be based on comparing the value at hand with the values obtained for the data distribution $\mathscr{D}$, which can be estimated, in practice, using a training set $D$:
@@ 3.4 Creating more meaningful explanations
In practice, we expect to sample $N$ different explanations, and, thereby, $N$ different yet representative prototypes for each class $y_{i}\in Y$ (assuming $m=c$). Although presenting all these prototypes as part of the explanation might be impractical, different strategies can be considered to summarize or distill the results, for the sake of more meaningful and useful explanations. First, among the prototypes corresponding to the most likely class, the ones closest to the input at hand can be considered the best candidates to justify the prediction of the model. Furthermore, the closest prototypes corresponding to the remaining classes can be taken as counterarguments against the most likely class, which can be used, for instance, to identify and explain why an input can be considered ambiguous or challenging to classify. On the other hand, the farther prototypes corresponding to the most likely class can be used to exemplify alternative yet still valid representations for that class, which can provide useful complementary information, thus being valuable for knowledge-acquisition purposes.
4 Training procedure
@@
and $\tau_{1},\tau_{2},\tau_{3}\in\mathbb{R}^{+}$ weight the contribution of each term. Notice that $\mathscr{L}_{\text{NLL}}$ represents the classification (cross-entropy) loss, averaged for $N$ samples of prototypes, and $\mathscr{L}_{\text{REC}}$ the reconstruction loss. Finally, $\mathscr{L}_{\text{INT}}$ can be seen as an interpretability loss, as it will encourage the distribution over the prototypes to be similar to the distribution of the encoded data, which is a crucial requirement for interpretability Li et al. (2018). Based on this methodology, the entire architecture can be jointly optimized by means of conventional gradient descent approaches.
@@ 5 Experiments
For illustration purposes, the effectiveness and capabilities of the proposed model will be assessed on the MNIST dataset Lecun et al. (1998), while additional experiments on the Fashion-MNIST Xiao et al. (2017) and K-MNIST Clanuwat et al. (2018) datasets are reported in Appendix D and Appendix E, respectively. In all the experiments, the Euclidean distance will be used as the distance metric $\phi$, and a Gaussian distribution will be assumed for each $p_{\lambda_{i}}(\bm{r}_{i})$, $1\leq i\leq c$. The full implementation details are reported in Appendix B.
Learned latent spaces and distributions over the prototypes
For illustration, the results obtained for the case in which $l=2$ (i.e., a two-dimensional latent space is considered), are shown in Figure 4. The left figure shows the learned latent space, visualizing the latent representations assigned to a random set of test data (dots), as well as the learned distribution over the prototypes $p_{\lambda_{i}}(\bm{r}_{i})$, $1\leq i\leq c$ (contour lines). The right figure represents 15 random sets of prototypes $R$ sampled from the learned distributions, and decoded by the decoder $g$. As can be seen, even for a very constrained (low-dimensional) latent space, the model is capable of capturing and decoding prototypical and varied representations for each class. The model achieves a top-1 accuracy of $0.98$ and $0.96$ in the train and test set, respectively, computed for $N=30$ inferences. Further results for $l=5$ and $l=10$ are shown in Appendix C, where it can be observed that the increase in the dimensionality of the latent space makes it possible to capture prototype distributions with a higher degree of complexity in their structure, as well as to obtain more detailed decodings. Further results for the Fashion-MNIST and K-MNIST datasets are provided in Appendix D and Appendix E, respectively. In addition, to show that our approach can scale to a larger number of classes, Appendix F includes results on the E-MNIST dataset Cohen et al. (2017), an extension of MNIST which contains handwritten representations of numbers and letters and is composed of 47 classes.
@@
Figure 4: Results obtained with a Prob-PSENN with $l=2$ for the MNIST dataset. (a) Learned latent space, where the contours represent the prototype distributions $p_{\lambda_{i}}(\bm{r}_{i})$ and the dots represent encoded input samples. (b) 15 random sets of prototypes $R_{n}=(r_{1},\dots,r_{c})\text{ s.t. }r_{i}\sim p_{\lambda_{j}}(\bm{r}_{j}),1\!\leq\!i\!\leq\!c$, $151\!\leq\!n\!\leq\!$ 15.
Comparison between conventional PSENN and Prob-PSENN
First, to thoroughly compare the predictive performance of both approaches, the accuracies obtained for the three datasets considered are compared in Table 1. For the sake of completeness, the performance of PSENNs is computed for different number of prototypes, while preserving its original architecture, consisting of a Convolutional Autoencoder, a latent space of $l=40$ dimensions, and using elastic deformation as data augmentation. The performance of Prob-PSENNs has been evaluated for the following configurations, fully described in Appendix B: using a Multi-Layer Perceptron Autoencoder (MLP), a Convolutional Autoencoder (CNN), and including elastic deformation Simard et al. (2003) as data augmentation (AUG). As can be seen, Prob-PSENNs achieve a very competitive performance in all the datasets, with an accuracy gain of $2\%$ on the K-MNIST dataset and a loss of less than $1\%$ on MNIST and F-MNIST, while having as benefits a greater robustness, trustworthiness and enhanced explanatory capabilities, as we showcase in Figure 5. In particular, this figure compares, for the MNIST task, the explanation computed by a conventional PSENN ($m=15$) with the one computed by Prob-PSENN ($l=5$, MLP). More comparisons will be provided in the supplementary material. As can be seen, the prototypes sampled by Prob-PSENN exhibit considerably more diversity than the fixed prototypes learned by a conventional PSENN, which enables to find more tailored prototypes for the input at hand, as well as to provide alternative forms that the class might take. In addition, while PSENNs only return point distances to the prototypes, Prob-PSENNs provide a distribution over those distances, enabling to assess the uncertainty of the prediction, and, hence, representing a more robust and trustworthy approach.
Figure 5: Comparison between the explanation provided by a PSENN with fixed prototypes (left) and a Prob-PSENN (right). For simplicity, the four most likely classes are visualized for Prob-PSENN, considering, for each class, the four closest prototypes and the most distant ones.
Table 1: Predictive accuracies (mean and standard deviation for 10 runs) obtained by Prob-PSENN and PSENN for the MNIST, Fashion-MNIST and K-MNIST datasets. The results are shown in percentages.
Configuration MNIST F-MNIST K-MNIST
PSENN ($l=40$, $m=10$) 98.97 $\pm$ 0.16 90.15 $\pm$ 0.33 91.11 $\pm$ 0.47
PSENN ($l=40$, $m=15$) 98.99 $\pm$ 0.11 90.10 $\pm$ 0.61 90.93 $\pm$ 0.41
PSENN ($l=40$, $m=20$) 99.01 $\pm$ 0.12 90.22 $\pm$ 0.50 91.15 $\pm$ 0.53
PSENN ($l=40$, $m=30$) 99.05 $\pm$ 0.15 89.90 $\pm$ 0.33 91.04 $\pm$ 0.96
PSENN ($l=40$, $m=40$) 98.93 $\pm$ 0.17 90.17 $\pm$ 0.39 90.81 $\pm$ 0.96
PSENN ($l=40$, $m=50$) 99.02 $\pm$ 0.11 89.76 $\pm$ 0.44 91.17 $\pm$ 0.74
Prob-PSENN ($l=5$,   MLP) 97.55 $\pm$ 0.29 88.83 $\pm$ 0.25 88.69 $\pm$ 1.23
Prob-PSENN ($l=10$, MLP) 97.59 $\pm$ 0.48 88.77 $\pm$ 0.50 88.92 $\pm$ 1.00
Prob-PSENN ($l=10$, CNN) 98.28 $\pm$ 0.17 89.47 $\pm$ 0.34 89.06 $\pm$ 0.34
Prob-PSENN ($l=10$, MLP + AUG) 98.53 $\pm$ 0.08 87.78 $\pm$ 0.50 93.17 $\pm$ 0.33
Prob-PSENN ($l=10$, CNN + AUG) 98.75 $\pm$ 0.17 87.98 $\pm$ 0.46 89.40 $\pm$ 1.14
Evaluating the explanatory uncertainties
@@
Figure 6 compares, for a Prob-PSENN with $l=2$, the areas of the latent space with high explanatory uncertainty. The aleatoric and epistemic explanatory uncertainties (second and third columns) have been evaluated using the metrics $\mathscr{U}_{A}(x)$ and $\mathscr{U}_{E}(x)$ defined in Equations (3) and (4), respectively. The fourth column shows the areas with high epistemic uncertainty exclusively (i.e., with no aleatoric uncertainty), which has been computed as $\mathscr{U}_{E}(x)-\mathscr{U}_{A}(x)$. As can be seen from the results, the regions with high aleatoric uncertainty are those that are at a similar distance from the two regions that the model associates to each class. On the other hand, the areas with high epistemic uncertainty are those far from the regions to which the prototype distributions assign a high density, and, therefore, the explanations are no longer representative of those regions.
@@
Figure 6: Evaluating the explanatory uncertainty in different regions of the learned latent space.
@@
Figure 7 compares the epistemic uncertainty achieved by a Prob-PSENN, trained for the MNIST dataset, when classifying 1000 test samples of the MNIST and Fashion-MNIST datasets. As can be seen, when inputs from the Fashion-MNIST dataset are presented to the model, a very high explanatory epistemic is captured for almost all the inputs, while for the MNIST test data the uncertainty is uniformly distributed. For reference, Figure 7 also shows, for each dataset, the samples that achieve the lowest and highest uncertainty. As can be seen, even for the MNIST dataset, highly unusual inputs that could be considered out-of-distribution instances can be found, which are effectively detected by the uncertainty estimation capabilities of Prob-PSENN.
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 7: Histograms showing the explanatory epistemic uncertainty obtained by a Prob-PSENN for 1000 test samples from MNIST (1st figure) and Fashion-MNIST (2nd figure). Notice that the scales of the Y-axis differ in the two plots. The following four images display those inputs that achieve, for each dataset, the lowest and highest epistemic uncertainty.
@@
Finally, the prediction and explanation for a particular input with high aleatoric uncertainty is shown in Figure 8. The results are shown for a Prob-PSENN with $l=5$ and $N=50$ inferences. The top row includes the density function over the distances corresponding to the two most likely classes, $\bm{\delta}_{x,4}$ and $\bm{\delta}_{x,9}$, estimated by Gaussian-kernel density estimation, as well as a box plot showing the variability in the output probabilities assigned to each class, considering all the inferences. As it can be assessed, the distributions of both the distances and the output probabilities are highly overlapped, evidencing a high explanatory and predictive aleatoric uncertainty. The second row of the figure includes the input and a subset of the sampled prototypes, sorted by distance to the input in increasing order. As can be seen, both distributions over the prototypes capture and sample prototypical representations that could be taken as equally descriptive of the input at hand, thus explaining its ambiguity.
@@
Harnessing uncertainty to improve predictive accuracy and reliability
@@
We experimentally validated that taking model uncertainty into consideration can provide substantial improvements in accuracy. Specifically, by discarding inputs with high explanatory epistemic uncertainty (e.g., greater than 95% or 99%), the accuracy of Prob-PSENN increases in all the datasets and configurations, as we show in Table 2 (Appendix G). To complement these results, Table 3 shows the percentage of inputs that are discarded in each case. At the same time, a large number of the discarded inputs are anomalous or challenging to classify, which would justify discarding them for the sake of higher reliability and robustness. As evidence, we show in Table 4 that the error rate of the deterministic baseline model is also higher in the inputs which receive high uncertainty than in the rest. Thus, this evaluation shows how capturing model uncertainty provides greater accuracy, robustness and safety. We justify the choice of the mentioned uncertainty metric, based on the fact that it is reasonable to consider a prediction unreliable if all the prototypes have low similarity to the input, given the distance-based classification pipeline of the model.

### v2

@@ 3.3.2 Quantifying explanation uncertainty
Regarding epistemic explanatory uncertainty, as explained above, high uncertainty will be achieved for those inputs for which the distances to all the prototypes are large. On that account, the minimum sample mean: $\bar{\delta}_{x,i*}=\min_{1\leq j\leq c}\bar{\delta}_{x,j}$ represents a suitable base metric for evaluating this uncertainty. Furthermore, in order to determine the severity of the uncertainty in a normalized and scale-independent manner, a quantile-based metric will be defined, close to the approach proposed in [33]. This metric will be based on comparing the value at hand with the values obtained for the data distribution $\mathscr{D}$, which can be estimated, in practice, using a training set $D$:
@@ 3.4 Creating more meaningful explanations
In practice, we expect to sample $N$ different explanations, and, thereby, $N$ different yet representative prototypes for each class $y_{i}\in Y$ (assuming $m=c$). Although presenting all these prototypes as part of the explanation might be impractical, different strategies can be considered to summarize or distill the results, for the sake of more meaningful and useful explanations. First, among the prototypes corresponding to the most likely class, the ones closest to the input at hand can be considered the best candidates to justify the prediction of the model. Furthermore, the closest prototypes corresponding to the remaining classes can be taken as counterarguments against the most likely class, which can be used, for instance, to identify and explain why an input can be considered ambiguous or challenging to classify. On the other hand, the farther prototypes corresponding to the most likely class can be used to exemplify alternative yet still valid representations for that class, which can provide useful complementary information, and is thus valuable for knowledge-acquisition purposes.
3.5 Training procedure
@@
and $\tau_{1},\tau_{2},\tau_{3}\in\mathbb{R}^{+}$ weight the contribution of each term. Notice that $\mathscr{L}_{\text{NLL}}$ represents the classification (cross-entropy) loss, averaged for $N$ samples of prototypes, $\mathscr{L}_{\text{REC}}$ the reconstruction loss, and $\mathscr{L}_{\text{INT}}$ an interpretability loss, as it will encourage the distribution over the prototypes to be similar to the distribution of the encoded data, which is a crucial requirement for interpretability [27]. Based on this methodology, the entire architecture can be jointly optimized by means of conventional gradient descent approaches.
@@ 4 Experiments
In this section, we describe the experimental setup designed to evaluate the effectiveness and capabilities of the proposed model, followed by a presentation and discussion of the corresponding results. The primary goal of these experiments is to demonstrate how Prob-PSENNs can: (i) learn highly diverse and prototypical representations of the classes, (ii) provide reliable uncertainty quantification for both the explanation and the classification, and (iii) leverage the uncertainty quantification mechanism to improve the accuracy, reliability and trustworthiness of the model.
4.1 Setup
The effectiveness and capabilities of the proposed model will be assessed on the MNIST [26], Fashion-MNIST [44], K-MNIST [8] and SVHN [31] datasets. This choice of datasets is based on the fact that they provide a clear and coherent notion of prototypes to facilitate the assessment of our results, particularly regarding explanation quality (e.g., prototype diversity) and uncertainty quantification (e.g., ambiguity).
The autoencoder module of Prob-PSENN will be implemented by a CNN-based Variational AutoEncoder (VAE), which adds an additional regularization term to the overall loss function, namely, the Kullback–Leibler (KL) divergence between the learned latent distribution and a prior (Gaussian) distribution [23]. The full implementation details are reported in Appendix B. For each prototype distribution $p_{\lambda_{i}}(\bm{r}_{i})$, $1\leq i\leq c$, a full covariance Gaussian distribution will be assumed, and, as in [27], the squared Euclidean distance will be used as the distance metric $\phi$ to measure the distance between the encoded input and the sampled prototypes (in the latent space). The final classification layer $W$ will be set as $W=-I$, where $I$ represents the identity matrix of dimension $c$, thus ensuring that the association between the prototype $\bm{r}_{i}$ and the class $y_{i}$, $1\leq i\leq c$, is explicit and unequivocal, further promoting transparency. The Adam optimization method [22] has been used to train the networks, using a learning rate of $0.001$. Our code is publicly available at https://github.com/vadel/Prob-PSENN.
4.2 Learned latent spaces and distributions over the prototypes
First, for illustration, the results obtained for the case in which $l=2$ (i.e., a two-dimensional latent space is considered) are shown in Figure 4, for the MNIST dataset. The left subfigure shows the learned latent space, visualizing the latent representations assigned to a random set of test data (dots), as well as the learned distribution over the prototypes $p_{\lambda_{i}}(\bm{r}_{i})$, $1\leq i\leq c$ (contour lines). The right subfigure represents 15 random sets of prototypes $R$ sampled from the learnt distributions, and decoded by the decoder $g$. As can be seen, even for a very constrained (low-dimensional) latent space, the model is capable of capturing and decoding prototypical and varied representations for each class.
@@ 4.2 Learned latent spaces and distributions over the prototypes
Figure 4: Illustrative results obtained with a Prob-PSENN with $l=2$ for the MNIST dataset. (a) Learned latent space, where the contours represent the prototype distributions $p_{\lambda_{i}}(\bm{r}_{i})$ and the dots represent encoded input samples. (b) 15 random sets of prototypes $R_{n}=(r_{1},\dots,r_{c})\text{ s.t. }r_{i}\sim p_{\lambda_{j}}(\bm{r}_{j}),1\!\leq\!i\!\leq\!c$, $151\!\leq\!n\!\leq\!$ 15.
Figure 5 shows randomly sampled sets of prototypes for the four datasets considered, for Prob-PSENNs trained with a latent dimensionality of $l=40$. As can be observed, the increase in the dimensionality of the latent space makes it possible to capture prototype distributions with a higher degree of complexity in their structure, and to obtain more detailed decodings. Furthermore, even for complex datasets such as K-MNIST (see Appendix E), the learned prototype distributions are able to capture representative, high-quality and diverse characterizations of the classes. More results are shown in Appendices C, D and E.
Figure 5: Visualization of prototypes sampled from the distributions learned with Prob-PSENNs ($l=40$), for the MNIST, F-MNIST, K-MNIST and SVHN datasets.
4.3 Evaluating and illustrating the explanatory uncertainties
@@
Figure 6 compares, for a Prob-PSENN with $l=2$ and the MNIST dataset, the areas of the latent space with high explanatory uncertainty. The aleatoric and epistemic explanatory uncertainties (second and third columns) have been evaluated using the metrics $\mathscr{U}_{A}(x)$ and $\mathscr{U}_{E}(x)$ defined in Equations (3) and (4), respectively. The fourth column shows the areas solely with high epistemic uncertainty (i.e., with no aleatoric uncertainty), which has been computed as $\mathscr{U}_{E}(x)-\mathscr{U}_{A}(x)$. As can be seen from the results, regions with high aleatoric uncertainty correspond to areas in the latent space that lie between distinct class clusters. On the other hand, the areas with high epistemic uncertainty are those far from the regions to which the prototype distributions assign high density, and, therefore, the explanations are no longer representative of those regions.
@@
Figure 6: Evaluating the explanatory uncertainty in different regions of the latent space learned by Prob-PSENN.
Figure 7: Histograms showing the explanatory epistemic uncertainty obtained by a Prob-PSENN (trained on the MNIST dataset) for 10000 test samples from MNIST, F-MNIST and K-MNIST. Notice that the scales of the Y-axis differ in the three plots.
@@
Figure 7 compares the epistemic uncertainty achieved by a Prob-PSENN, trained for the MNIST dataset, when classifying 1000 test samples of the MNIST, F-MNIST and K-MNIST datasets. As can be seen, when inputs from the F-MNIST and K-MNIST dataset are presented to the model, a very high explanatory epistemic is captured for almost all the inputs, while for the MNIST test data the uncertainty is uniformly distributed. Furthermore, Figure 8 shows examples of MNIST inputs that achieve the lowest and highest uncertainties. As can be seen, even for the MNIST dataset, highly unusual inputs that could be considered out-of-distribution instances can be found, which are effectively detected by the uncertainty estimation capabilities of Prob-PSENN.
(a) High epistemic explanatory uncertainty.
(b) Low epistemic explanatory uncertainty.
Figure 8: Examples of inputs receiving (a) high or (b) low explanatory epistemic uncertainty from a Prob-PSENN trained on the MNIST dataset.
@@
Finally, the prediction and explanation for a particular input with high aleatoric uncertainty is shown in Figure 9. The top row includes the density function over the distances corresponding to the two most likely classes, $\bm{\delta}_{x,4}$ and $\bm{\delta}_{x,9}$, estimated by Gaussian-kernel density estimation, as well as a box plot showing the variability in the output probabilities assigned to each class, considering all the inferences. Note that the distributions of both the distances and the output probabilities are highly overlapping, evidencing a high explanatory and predictive aleatoric uncertainty. The second row of the figure includes the input and a subset of the sampled prototypes, sorted by distance to the input in increasing order. As can be seen, both distributions over the prototypes capture and sample prototypical representations that could be taken as equally descriptive of the input at hand, thus explaining its ambiguity.
@@
4.4 Comparison between conventional PSENN and Prob-PSENN
To analyze the predictive performance of Prob-PSENNs, we begin by comparing the accuracies obtained for the MNIST, F-MNIST and K-MNIST test sets, as shown in Table 1. For reference, we include the results of the deterministic baseline models proposed in Li et al. [27] and Gautam et al. [13], reproduced using the implementation released by the authors (see Appendix A for a further discussion on these works). The results show that Prob-PSENNs achieve either superior or competitive performance in all the datasets, while having as benefits greater robustness, trustworthiness, and enhanced explanatory capabilities. These capabilities are showcased in Figure 10, which compares, for the MNIST task, the explanation computed by a conventional PSENN ($m=15$) with that computed by Prob-PSENN. More comparisons are provided in Appendices D and E. As can be seen, the prototypes sampled by Prob-PSENN exhibit considerably more diversity than the fixed prototypes learned by a conventional PSENN, which enables one to find more tailored prototypes for the input at hand, as well as to provide alternative forms that the class might take. In addition, while PSENNs only return point distances to the prototypes, Prob-PSENNs provide a distribution over those distances, enabling assessment of the uncertainty of the prediction, and, hence, accounting for a more trustworthy approach.
Table 1: Predictive accuracies (mean and standard deviation for 5 runs) obtained for the MNIST, Fashion-MNIST and K-MNIST datasets. The accuracies are shown in percentages. Best results in the default setting are in bold, while the best results in the uncertainty-aware setting (Unc. Aware; see Section 4.4.1) are in bold and italics.
Model Configuration MNIST F-MNIST K-MNIST
Li et al. [27] ($m=10,l=40$) 99.0 $\pm$ 0.16 90.2 $\pm$ 0.33 91.1 $\pm$ 0.47
Li et al. [27] ($m=15,l=40$) 90.0 $\pm$ 0.11 90.1 $\pm$ 0.61 90.9 $\pm$ 0.41
Li et al. [27] ($m=30,l=40$) 99.1 $\pm$ 0.15 89.9 $\pm$ 0.33 91.0 $\pm$ 0.96
Li et al. [27] ($m=50,l=40$) 99.0 $\pm$ 0.11 89.8 $\pm$ 0.44 91.2 $\pm$ 0.74
Gautam et al. [13] ($m=50,l=256$) 99.4 $\pm$ 0.10 91.9 $\pm$ 0.20 96.1 $\pm$ 0.16
Prob-PSENN ($m\!=\!10$, $l\!=\!20$) 99.4 $\pm$ 0.05 91.9 $\pm$ 0.14 96.7 $\pm$ 0.21
Prob-PSENN ($m\!=\!10$, $l\!=\!40$) 99.4 $\pm$ 0.02 92.0 $\pm$ 0.24 96.6 $\pm$ 0.17
     + Unc. Aware ($\alpha=0.99$) 99.9 $\pm$ 0.02 95.6 $\pm$ 0.35 99.5 $\pm$ 0.03
     + Unc. Aware ($\alpha=0.95$) 99.9 $\pm$ 0.01 97.3 $\pm$ 0.21 99.7 $\pm$ 0.02
Figure 10: Comparison between the explanation provided by a PSENN [27] with fixed prototypes (left) and a Prob-PSENN (right). For simplicity, the results corresponding to the four most likely classes are visualized for Prob-PSENN. For each of these classes, we display the four prototypes closest to the input, as well as the most distant prototype.
4.4.1 Harnessing uncertainty to improve predictive accuracy and reliability
@@
We experimentally validated that taking model uncertainty into consideration can provide substantial improvements in accuracy. Specifically, by discarding inputs with high explanatory epistemic uncertainty [^1] We justify the choice of the mentioned uncertainty metric based on the fact that it is reasonable to consider a prediction unreliable if all the prototypes have low similarity to the input, given the distance-based classification pipeline of the model. (e.g., greater than $\alpha=95$% or $\alpha=99$%), the accuracy of Prob-PSENN increases in all the datasets and configurations, as we show in Table 1 (last two rows). To complement these results, Table 2 shows the percentage of inputs that are discarded in each case. The trade-off between the accuracy and the ratio of discarded samples (i.e., those with highest epistemic uncertainty) is also shown in Figure 11. At the same time, a large number of the discarded inputs are anomalous or challenging to classify, which would justify discarding them for the sake of higher reliability. As evidence, we show in Table 2 that the error rate of the deterministic baseline model is dramatically higher in the inputs which receive high uncertainty than in the rest. For instance, the average error rate percentage of the model in the discarded inputs ($\alpha=0.99$) is roughly 23%, 40% and 51% for MNIST, F-MNIST and K-MINST, respectively. Thus, this evaluation shows how capturing model uncertainty provides greater accuracy, reliability and safety.
Figure 11: Trade-off between the accuracy gains and the discarded ratio, when the inputs with highest epistemic uncertainty are rejected. Results averaged for 5 runs. Notice that the scales of the Y-axis differ in the three plots.
Table 2: Percentage of inputs discarded by Prob-PSENNs at different uncertainty thresholds ($\alpha$) and the error rates (%) of the deterministic baseline model [27] for high-uncertainty (discarded) and low-uncertainty (retained) inputs across MNIST, F-MNIST, and K-MNIST.
Threshold Metric MNIST F-MNIST K-MNIST
$\alpha=0.999$ Discarded Inputs   1.38 $\pm$ 0.10   2.23 $\pm$ 0.49   5.84 $\pm$ 0.21
Error Rate (High-Unc.) 31.30 $\pm$ 3.22 45.49 $\pm$ 2.33 64.20 $\pm$ 2.61
Error Rate (Low-Unc.)    0.79 $\pm$ 0.20    9.05 $\pm$ 0.57    6.28 $\pm$ 0.98
$\alpha=0.99$ Discarded Inputs   3.02 $\pm$ 0.03   9.19 $\pm$ 0.73 11.22 $\pm$ 0.23
Error Rate (High-Unc.) 23.72 $\pm$ 2.86 40.01 $\pm$ 1.73 50.95 $\pm$ 2.73
Error Rate (Low-Unc.)    0.51 $\pm$ 0.18    6.80 $\pm$ 0.46    4.44 $\pm$ 0.85
$\alpha=0.95$ Discarded Inputs   6.78 $\pm$ 0.08 17.52 $\pm$ 0.70 19.84 $\pm$ 0.28
Error Rate (High-Unc.) 13.15 $\pm$ 1.68 33.28 $\pm$ 1.56 36.59 $\pm$ 2.75
Error Rate (Low-Unc.)    0.35 $\pm$ 0.16    4.87 $\pm$ 0.34    3.00 $\pm$ 0.63
4.4.2 Ablation studies
To isolate the impact of the probabilistic reformulation of the prototypes, Appendix F compares the performance of Prob-PSENN with its deterministic counterpart, using the same architecture and training pipeline. The point-wise prototypes have been optimized using the loss function proposed in [27]. The results obtained with deterministic PSENNs are summarized in Table F.1, computed considering $l\!\in\!\{20,40\}$ latent dimensions and $m\!\in\!\{10,30\}$ prototypes. The results for Prob-PSENN are included in the last rows of the table, as reference. As can be observed, both approaches achieve nearly identical performance, with average accuracy differences below 0.2%, further demonstrating that the advantages of our probabilistic model do not imply a reduction in its predictive performance.
In addition, Appendix F further examines the benefits of using a VAE over a vanilla AE as the backbone architecture. Table F.2 compares the predictive performance of Prob-PSENNs when replacing the VAE with an AE while keeping all other architectural and training details unchanged. The results show that AEs lead to a performance drop ranging from approximately 0.1% to 1.2%, depending on the dataset. In addition, Figure F.21 visualizes randomly sampled prototypes from Prob-PSENNs with different backbone architectures, showing that while vanilla AE backbones provide comparable results in lower latent dimensions (e.g., $l=10$), they struggle to maintain high-quality prototype distributions as the dimensionality increases. In contrast, VAEs enable effective scaling to higher-dimensional spaces. This demonstrates that, while Prob-PSENNs are not inherently restricted to VAE backbones, more regularized latent spaces enhance both predictive performance and prototype quality in high-dimensional settings.
4.5 Scalability
Table 3: Results on the SVHN dataset.
Configuration Accuracy (%)
Gautam et al. [13] (ProtoPNet) $88.6\pm 0.3$
Gautam et al. [13] (ProtoVAE) $92.2\pm 0.3$
Gautam et al. [13] (SENN) $91.5\pm 0.4$
Prob-PSENN $92.1\pm 0.2$
In order to show that our approach can scale to larger numbers of classes, Appendix G includes results on the E-MNIST dataset [9], an extension of MNIST which contains handwritten representations of numbers and letters and is composed of 47 classes. As can be seen, Prob-PSENN achieves a highly diverse and representative sets of prototypes for all the classes, which would require an excessively large number of pointwise prototypes to achieve using deterministic approaches. Furthermore, to show that Prob-PSENNs can scale to colour datasets, we tested our model on the SVHN dataset. The accuracy of the model in this dataset is reported in Table 3. The results show that the predictive performance of Prob-PSENNs is on par with previous deterministic approaches, when not superior.